paul’s demand function for good x is given by x(px,ly,m)=2m/5px. his income, m, is £1000, the price of x,px, is £5 and the price of y, pay, is £20.
what will be the change in paul’s demand for good x if the price of good falls to £4?
1
Expert's answer
2020-11-17T12:39:57-0500

Demand: "\frac{2m}{5px}"

Income: £1000

Price (Px): £5

Price (Py): £20

Demand = "(2\times1000)/(5\times5) = 80"

If Px = £4

Demand = "(2\times1000)/(5\times4) = 100"

Change = "\frac{(100-80)}{80} = 25\%"